[Hallc_running] [New Logentry] RC Daily Update

dhamilto at jlab.org dhamilto at jlab.org
Wed Dec 5 10:00:03 EST 2018

Logentry Text:

- Production data-taking continues with the SHMS at 9.79 degrees. Swing was busy, with an arc energy measurement, HMS optics run and dummy target data-taking. Back to production for Owl.
- Mark is still looking into the issue of lower HMS tracking efficiency for singles events.

RC Meeting:
- There will be an RC meeting today with Carlos at 4:15pm, second floor counting house.

Notices in effect for shift crews:
- Shift leaders, please fill out the check list once per shift.
- Shift leaders, please do a quick cross-check of the BTA numbers against your shift summary.
- Check that the pre-scaler gui is active and hasn't frozen / doesn't need re-started.
- Be on the lookout for any errors when running the replay, particularly errors concerning missing reference times.
- Note the new beam positions for this energy (on the whiteboard / note on wall)
- NB If the HV trips for SHMS S1Y5B contact Simona M. '''DO NOT change the HV from the nominal value!'''

Planned accelerator configuration changes:
- Energy change on Monday, 6am (note the updated time).


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3635172
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181205/6f13dbeb/attachment.html>

More information about the Hallc_running mailing list